Numerical Analysis I, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 09

1. Section 4.3: Problems 7, 14, 15, 16, 18, 19, 20.

2. Section 4.4: Problems 7, 14, 16, 17, 19, 24. For problem 7, write your code directly without using the sample codes.